Given the equation p = q2 + 5q +2. Find the values of q corresponding to
p = -2.

Found 2 solutions by nerdybill, mananth:
Answer by nerdybill(7384)

Given the equation p = q2 + 5q +2. Find the values of q corresponding to p = -2
.
-2 = q^2 + 5q + 2
0 = q^2 + 5q + 4
0 = (q+4)(q+1)
q = {-4, -1}

Answer by mananth(16949)

p = q2 + 5q +2. Find the values of q corresponding to
p = -2.
..
-2=q^2+5q+2
add +2 to both sides
q^2+5q+4 =0
q^2+4q+q+4=0
q(q+4)+1(q+4)=0
(q+1)(q+4)=0
q=-1 OR q=-4